OpenStudy (camzzzie):

PLEASE HELP -2x+3(9+2x)-5=-10

OpenStudy (camzzzie):

9+2x was y so i subbed it i but i dont know how to continue it

OpenStudy (anonymous):

Simplifying 2x + -10 = 10 + -3x Reorder the terms: -10 + 2x = 10 + -3x Solving -10 + 2x = 10 + -3x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '3x' to each side of the equation. -10 + 2x + 3x = 10 + -3x + 3x Combine like terms: 2x + 3x = 5x -10 + 5x = 10 + -3x + 3x Combine like terms: -3x + 3x = 0 -10 + 5x = 10 + 0 -10 + 5x = 10 Add '10' to each side of the equation. -10 + 10 + 5x = 10 + 10 Combine like terms: -10 + 10 = 0 0 + 5x = 10 + 10 5x = 10 + 10 Combine like terms: 10 + 10 = 20 5x = 20 Divide each side by '5'. x = 4 Simplifying x = 4

OpenStudy (camzzzie):

thank you but 4 isnt one of the multiple choice options:(

OpenStudy (anonymous):

damn

OpenStudy (camzzzie):

I'm gonna type out the original problem anfd see if you can tell me what i did wrong

OpenStudy (anonymous):

-2x+3(9+2x)-5=-10 -2x + 27+ 6x -5 = -10 4x + 22 = -10 -22 -22 4x = -32 -- --- 4 4 x = -8

OpenStudy (camzzzie):

2x-y+z=-4 z=5 -2x+3y-z=-10 solve by substitution

OpenStudy (anonymous):

oh i messed up

OpenStudy (camzzzie):

thats ok. thank you both for your help!

OpenStudy (anonymous):

but camzzzie i gave you a medal

OpenStudy (camzzzie):

i just gave you one on some other question so we good?:)

OpenStudy (anonymous):

do you understand how i got it?